# Insecurity Through Obscurity: Veiled Vulnerabilities in Closed-Source Contracts

### v1

@@ Abstract.
Most blockchains cannot hide the binary code of programs (i.e., smart contracts) running on them. To conceal proprietary business logic and to potentially deter attacks, many smart contracts are closed-source and employ layers of obfuscation. However, we demonstrate that such obfuscation can obscure critical vulnerabilities rather than enhance security, a phenomenon we term insecurity through obscurity. To systematically analyze these risks on a large scale, we present skanf, a novel EVM bytecode analysis tool tailored for closed-source and obfuscated contracts. skanf combines control-flow deobfuscation, symbolic execution, and concolic execution based on historical transactions to identify and exploit asset management vulnerabilities. Our evaluation on real-world Maximal Extractable Value (MEV) bots reveals that skanf detects vulnerabilities in 1,028 contracts and successfully generates exploits for 373 of them, with potential losses exceeding \$9.0M. Additionally, we uncover 40 real-world MEV bot attacks that collectively resulted in \$900K in losses.
1. Introduction
Smart contracts (Wood et al., 2014) are autonomous programs on a blockchain, offering strong integrity and availability guarantees. However, because most blockchains lack privacy, smart contract bytecode is publicly visible. While transparency is a key feature, it becomes a drawback when a smart contract’s logic embodies competitive advantages or proprietary strategies.
To mitigate this, a variety of obfuscation techniques are commonly applied. A prominent example is the Maximal Extractable Value (MEV) bots, which are optimized smart contracts deployed by MEV searchers to perform arbitrary, sandwich attacks and other on-chain actions to extract MEV (Daian et al., 2020). Six of the top 10 MEV bots, ranked by the number of transaction bundles they sent (libMEV, 2025), employ some form of obfuscation.
@@ 1. Introduction
There are several plausible reasons to use obfuscation, including to keep proprietary strategies private and to deter attacks (i.e., security from obscurity). However, our analysis of the recent “Destroyer Inu” incident (detailed in Sec. 3) reveals that obfuscation can also introduce insecurity. In this case (detailed in Sec. 3), 22 ETH (worth \$51,056 at the time) were stolen from a closed-source [^1] A smart contract is closed-source if its source (e.g., Solidity code) is not publicly available. Binary code is always publicly available on Ethereum. smart contract equipped with layers of obfuscation. The attack exploited a well-known vulnerability—tx.origin phishing (by Example, 2025)—that standard analysis tools can readily detect. But because the contract was obfuscated, the vulnerability went unnoticed until it was exploited. In this case, obscurity arguably renders the contract less secure by hindering analysis, making this a prime instance of insecurity through obscurity (Meunier, 2008) [^2] Compare this with modern cryptographic schemes following Kerckhoffs’s principle that a system’s security should not be compromised if adversaries uncover its method of operation (Petitcolas, 2019), or, in the words Shannon: “the enemy knows the system” (Shannon, 1949)..
@@
This work.   Due to the lack of effective analysis tools, little is known about the security of closed-source smart contracts, especially those that handle large volumes of assets and present lucrative targets for attackers. In this paper, we aim to develop effective code analysis techniques to bridge this gap.
We focus on a broad class of security vulnerabilities that we refer to as asset management vulnerabilities, issues that allow an adversary to manipulate a smart contract’s logic to steal assets (i.e., transferring tokens from the victim to the attacker herself). While our techniques are general, we use MEV bots as a real-world testbed to evaluate the efficacy of our tool and as concrete examples to study the security of closed-source smart contracts in the wild.
Our goals can be summarized as the following three research questions (RQs):
@@
  • RQ2: How to effectively detect asset management vulnerabilities in closed-source smart contracts? Even if obfuscation is removed, detecting such vulnerabilities can be challenging, given the complexity of MEV bots. Existing tools often timeout when applied to MEV bots, even without obfuscation.
@@
We answer these questions using the following methods. First, we analyze notable attacks on closed-source contracts, and distill the vulnerabilities that allowed them and the methods used to execute them. Then, we use our new-found understanding to devise a novel deobfuscator that successfully uncovers crucial parts of closed-source contracts which are opaque to established tools. Making use of our deobfuscator and insights, we build a vulnerability inspector and an automatic attack generator. We combine our deobfuscator, vulnerability inspector and attack generator into a tool we call skanf  [^3] The name skanf stands for Sen, Kaihua, Aviv ’n Fan.. By applying skanf to closed-source contracts we collect, we count the number of closed-source contracts that have potential security vulnerabilities and automatically synthesize the exploits against them. We further estimate a lower bound of the amount of funds that can be taken from them using historical Ethereum data.
1.1. Challenges
@@
Numerous tools (Dedaub, 2025; ConsenSys, 2025; Frank et al., 2020; Gritti et al., 2023; Grech et al., 2019) are available to identify and exploit smart contract bugs. To illustrate the challenges in answering our research questions, we highlight three core limitations of existing tools that render them inefficient for our tasks.
Control-flow obfuscation.   Mainstream languages like Solidity and Vyper rely on jump tables with static jump destinations to manage function dispatch and control flow. However, contracts may employ “indirect jumps” to make the control flow dependent on runtime values to achieve obfuscation.
To the best of our knowledge, no existing tools effectively handle control flow obfuscation. State-of-the-art static analysis tools, such as Gigahorse (Grech et al., 2019, 2022; Lagouvardos et al., 2024), classify basic blocks as unreachable when a contract relies on indirect jumps to determine control flow. Decompilation tools such as Dedaub, which claims to “successfully decompile over 99.98% of deployed contracts on the Ethereum blockchain” (Dedaub, 2025), fail to function effectively when facing control flow obfuscation. For instance, Dedaub can only analyze 4% of code in the top-2 obfuscated MEV bots. Symbolic execution tools like Mythril (ConsenSys, 2025), ETHBMC (Frank et al., 2020), and Greed (Gritti et al., 2023; Ruaro et al., 2024) raise errors when jump destinations are symbolic.
@@
Lack of fine-grained analysis.   Deciding asset management vulnerabilities requires a fine-grained analysis of a smart contract’s logic, and simple pattern matching as done by existing tools does not suffice. For example, JACKAL (Gritti et al., 2023) flags CALL instructions with a fixed function selector as safe, which overlooks vulnerabilities where the attacker manipulates function parameters (instead of the function selector). As another example, one might flag tx.origin-based authentication as insecure (because it appears similar to tx.origin phishing (by Example, 2025)), but later steps may rectify the problem.
Performance due to complex logic.   In addition to asset management, the contracts we focus on also involve trading, flash loans, swaps on different DEXs, and other operations and interactions. This complexity leads to performance issues in static analysis, as the tools must analyze many control-flow paths and resolve intricate data-flow dependencies, increasing both processing time and memory usage. Moreover, the rapid growth of execution paths and the complexity of constraints make symbolic execution inefficient due to path explosion and the high computational cost of constraint solving. While some tools support concolic execution (ConsenSys, 2025), generating high-quality seed input remains a manual process that does not scale.
1.2. Our Methods
Recovering control flow.   We propose a method to de-obfuscate control flows exploiting the following EVM feature: Unlike traditional instruction sets such as x86-64 or ARM64, where jump instructions can target any executable address, the EVM requires all valid jump destinations to be explicitly marked with a JUMPDEST instruction (Wood et al., 2014). This allows us to determine the boundaries of basic blocks directly from the bytecode and reconstruct jump tables, even when jump destinations are determined dynamically at runtime. Our tool can instrument the bytecode with a reconstructed jump table to significantly improve the code coverage of subsequent analysis, which is crucial for efficacy.
Detecting asset management vulnerabilities.   To not rely on a specific pattern for detecting asset management vulnerabilities, we broadly look for vulnerable calls in the bytecode of a smart contract that an adversary can trigger with adversarial parameters. As a naive example, if a smart contract exposes a CALL instruction where the adversary can set the target function as well as its input, the adversary can call WETH.transfer with the “to” parameter set to itself to drain the victim’s WETH. Many attacks, however, do not require the adversary to control the entirety of the inputs.
Our approach draws inspiration from traditional program vulnerability detection, where the combination of symbolic execution and taint analysis has been used to address similar problems (Schwartz et al., 2010). To detect asset management vulnerability, we first use symbolic execution to identify execution paths leading to a CALL statement. Then, we apply taint analysis to examine whether the inputs to the CALL statement can be controlled by an adversary. Attacks are possible even if the adversary cannot control the entirety of the inputs, so we carefully enumerate cases where an adversary can manipulate critical parameters, such as the recipient address for approve or transfer.
Concolic execution based on historical transactions.   As mentioned above, a naive application of symbolic execution does not scale to handle real-world smart contracts of interest. Concolic execution solves the performance issues of symbolic execution by combining concrete and symbolic execution to mitigate path explosion (Sen, 2007). By leveraging concrete seed inputs to guide path exploration and simplify constraint solving, concolic execution can reduce the search space, improving efficiency. The “quality” of the seed input is critical, but there are no general algorithms to generate high-quality ones.
We observe, however, that high-quality inputs are already readily available on the blockchain: all historical transactions are publicly visible, and many of them invoke CALL instructions through various paths. This allows us to automatically extract high-quality seed inputs from external and internal transactions recorded on the blockchain, significantly improving runtime performance and the efficacy of vulnerability detection.
1.3. Implementation and Evaluation of skanf
Putting these ideas together, we developed skanf based on Gigahorse (Grech et al., 2019) and Greed (Gritti et al., 2023; Ruaro et al., 2024). At a high level, skanf takes the bytecode and the historical transactions of a given smart contract as the input and outputs exploits if the smart contract is vulnerable.
The workflow of skanf consists of three stages. First, it performs static analysis to detect indirect jumps and attempts to deobfuscate the control flow by inserting a switch table that converts all indirect jumps into direct jumps. Second, skanf parses the contract’s historical transactions as concrete seed input for concolic execution and identifies asset management vulnerabilities during the execution. For each identified vulnerability, skanf synthesizes exploits using constraint-solving, with adversary-controlled parameters — such as the ERC-20 token address and transfer amount — as constraints. The exploit is then validated and reported.
Evaluation results.   To evaluate the efficacy of skanf against real-world smart contracts, we compile a dataset of 6,554 MEV bot contracts and assess skanf alongside three state-of-the-art tools. For each tool, we limit the execution time to 10 minutes and count the number of vulnerable smart contracts it can detect. Our evaluation shows that skanf outperforms the other tools: it detects 1,028 vulnerable smart contracts and successfully generates exploits for 373 of them. If exploited, the potential losses of these vulnerabilities would exceed 9.0 million USD. In comparison, other tools such as ETHBMC (Frank et al., 2020) and JACKAL (Gritti et al., 2023) detected 0 and 18 vulnerable contracts, respectively. While Mythril identified 574 vulnerable contracts, our verification confirmed that 84% of them are false positives.
To understand if skanf can flag attacks that already happened, we conduct an empirical study on MEV bot attacks in the wild. We discovered 40 exploits targeting 12 MEV bot contracts, resulting in a total loss of \$900K. This confirms that asset management vulnerabilities pose a serious threat to bot security. All exploited contracts are flagged by skanf—had skanf been available to developers, these exploits could have been detected early on. Even after these attacks occurred, only three had been previously reported, probably because no tool was available to automatically scan for such exploits.
@@ 1.3. Implementation and Evaluation of skanf
In short, we make the following contributions:
@@
  • Building on these two techniques, we implement skanf. We plan to release the code publicly.
@@
  • We evaluate the efficacy of skanf against real-world closed-source and obfuscated smart contracts. skanf identifies 1,028 vulnerable contracts and automatically exploits 373 of them. The potential losses of these vulnerabilities exceed \$9.0M.
@@
  • We discover 40 real-world attacks targeting MEV bots, which we categorize as MEV phishing attacks. These attacks have resulted in an estimated loss of \$900K to searchers. All of the attacked smart contracts are flagged by skanf.
Responsible Disclosure.   As our exploits can readily compromise existing smart contracts, we responsibly disclosed them to the affected parties through a dedicated blockchain messaging service (Blockscan Chat by Etherscan (Etherscan Team, 2023)), following the same disclosure practice adopted by prior research (Qin et al., 2023). Moreover, we adhere to several additional responsible practices. We test all exploits on a local replication of the Ethereum blockchain that is disconnected from the public network. Unless the attacks have already occurred in the wild and the vulnerabilities are publicly known, we do not disclose the specific addresses of vulnerable contracts in the paper.
2. Preliminaries
2.1. Ethereum and Smart Contracts
@@
Ethereum.   Ethereum is the second-most popular blockchain, and the foremost smart contract-enabled blockchain. It relies on a Proof-of-Stake (PoS) mechanism, wherein actors can become part of the set of validators operating the blockchain by staking (i.e., depositing) Ethereum’s native token, ETH. Ethereum’s mechanism sequentially assign one validator to propose a block, i.e., a batch of transactions, every 12 seconds. The validator chosen to act as a proposer has a free hand in constructing its block.
@@
MEV supply chain.   Today’s Ethereum block construction pipeline involves several specialized entities. At one end, we have searchers who identify profitable opportunities, e.g., transactions that can be front-run (Zhou et al., 2023), and assemble transaction bundles that exploit them. These bundles are then sent to builders, who specialize in constructing profitable blocks, and who compete in an auction-esque process to have their blocks chosen by upcoming proposers.
@@
Smart contract.   Smart contracts are computer programs stored and executed on the blockchain. Once deployed, contracts can be invoked and interacted with using transactions. Ethereum contracts are specified in EVM bytecode, but can be written using high-level languages like Solidity (Team, 2025b) and then compiled for deployment (Wood et al., 2014). Besides using high-level languages, sophisticated developers can further optimize contracts–such as reducing gas costs–by directly writing bytecode or using low-level languages like Huff (Team, 2025a). While Ethereum smart contracts are stored on-chain using low-level EVM bytecode, developers can choose to open-source their contracts’ high-level code on platforms like Etherscan (Etherscan, 2025a).
The Application Binary Interface (ABI) specifies how to encode function calls and decode responses when interacting with a smart contract on Ethereum. If a contract is open-sourced, its ABI becomes publicly accessible; otherwise, the ABI is unavailable, making it difficult for the public to know how to interact with the contract.
The EVM is the execution environment for Ethereum smart contracts. Specifically, smart contracts execute in a stack-based virtual machine, meaning operations are performed by manipulating data on a last-in-first-out (LIFO) stack rather than via registers. Contracts are specified using low-level EVM bytecode composed of opcodes such as CALL, JUMP, and JUMPI. Control flow within the EVM is managed through the JUMP and JUMPI instructions, which provide unconditional and conditional branching, respectively.
@@
Indirect jump.   In the EVM, an indirect jump occurs when a jump instruction (JUMP/JUMPI) is executed and the top stack value (i.e., the jump destination) is not a statically known constant, but rather a runtime-computed value. Obfuscation techniques often rely on indirect jumps to hide control flow, making it difficult to statically determine which path may be executed.
@@
External call.   During execution, a smart contract can invoke another contract or externally owned account (EOA) using one of four opcodes: CALL, CALLCODE, DELEGATECALL, and STATICCALL. These opcodes enable the caller to pass along Ether (value), execution gas, and input data (calldata), facilitating inter-contract interactions. They differ in how the execution context is propagated, particularly with respect to msg.sender, tx.origin, and storage access. In this paper, we focus on the CALL instruction, which underlies common patterns such as calling an ERC-20 token’s transfer function; we refer to such invocations as external calls.
@@
We refer to the target contract that an external call aims to invoke as the target address. The calldata consists of two parts: the first four bytes, named function selector, which identify the target function to be executed in the target contract; and the remaining bytes, which specify the arguments passed to that function.
@@
Asset management.   In this paper, we focus on ERC-20 token assets held by smart contracts. The ERC-20 standard defines two methods for transferring tokens (Ethereum Foundation, 2025). An account can invoke transfer on an ERC-20 token contract to send tokens directly to another account. Alternatively, it can use approve to authorize another account to transfer tokens from its balance using transferFrom.
2.2. Threat Model

### v3

@@ Abstract
Most blockchains cannot hide the binary code of programs (i.e., smart contracts) running on them. To conceal proprietary business logic and to potentially deter attacks, many smart contracts are closed-source and in many cases exhibit code obfuscation, either intentionally introduced to hide internal logic or unintentionally produced by optimizations. However, we demonstrate that such obfuscation can obscure critical vulnerabilities rather than enhance security, a phenomenon known as insecurity through obscurity. To systematically analyze these risks on a large scale, we present skanf, a novel EVM bytecode analysis tool tailored for closed-source and obfuscated contracts. skanf combines control-flow deobfuscation, symbolic execution, and concolic execution based on historical transactions to identify and exploit asset management vulnerabilities. Our evaluation on real-world Maximal Extractable Value (MEV) bots reveals that skanf detects vulnerabilities in 1,030 contracts and successfully generates exploits for 394 of them, with potential losses of \$10.6M. Additionally, we uncover 104 real-world MEV bot attacks that collectively resulted in \$2.76M in losses.
1 Introduction
On public blockchains such as Ethereum, smart contract bytecode—and source code, if developers choose to release it—are publicly accessible. While transparency is a core feature, it allows a contract’s logic to be examined, reverse-engineered, or copied. As a result, many developers choose not to publish the source code of their smart contracts. Furthermore, bytecode obfuscation is commonly observed, either intentionally introduced by developers to protect their contracts from attacks [16, 17, 45], or as a byproduct of compiler optimizations in languages such as Vyper [69] and Huff [67].
To provide context, our study (App. A) of Ethereum shows that about 15% of the most active contracts (top 50K, ranked by the number of transactions interacted with them) are closed-source. Among these, obfuscated contracts collectively hold more than \$37M worth of crypto assets.
While obfuscation may have certain benefits due to “security through obscurity,” (e.g., some practitioners deem obfuscation necessary for deterring frontrunning attacks [17]), this paper is motivated by the observation that obfuscation can also introduce insecurity.
@@ 1 Introduction
To see the motivation, consider the recent “Destroyer Inu” incident, where 22 ETH (worth \$51,056 at the time) were stolen from a closed-source smart contract equipped with layers of obfuscation. The attack exploits a well-known vulnerable pattern (misuse of tx.origin [8] followed by improper asset management), but existing tools could not detect the full exploit because of control flow obfuscation. E.g., Mythril [12] recognizes the superficial misuse of tx.origin but misses the much more severe asset management vulnerability, only reporting a “low” risk. In this case, obscurity arguably renders the contract less secure by hindering analysis, making this a prime instance of insecurity through obscurity [47] [^1] Compare this with modern cryptographic schemes following Kerckhoffs’s principle that a system should not be compromised if adversaries uncover its method of operation [54], or, as Shannon said: “the enemy knows the system” [62]..
@@
This work.  Due to the lack of effective analysis tools, little is known about the security of closed-source smart contracts, particularly those handling large volumes of assets and are attractive to attackers. In this paper, we aim to develop effective code analysis techniques to address this gap. We focus on a broad class of security vulnerabilities that we refer to as asset management vulnerabilities. These vulnerabilities allow an adversary to steal assets, creating a strong incentive for exploitation if the vulnerabilities cannot be detected promptly.
Real-world testbed: MEV bots.  While our tool is general, we use MEV (Maximal Extractable Value) bots as the testbed for efficacy evaluation and as concrete examples to study the security of closed-source smart contracts in the wild.
MEV bots are highly optimized smart contracts deployed by MEV searchers to facilitate MEV extraction [14]. While some MEV bots are considered toxic, many play a beneficial role [20, 29]. For example, arbitrage bots align DEX prices with the broader market, providing users with accurate prices.
We choose MEV bots as our testbed because they are commonly closed source and obfuscated (we identify 6,554 closed-source MEV bots as of the time of writing; Six of the top 10, ranked by the number of transaction bundles they sent [42], exhibit some form of obfuscation.) Moreover, they handle large volumes of assets, making any asset management vulnerabilities particularly relevant and detrimental.
We summarize our goals as three research questions (RQs):
@@
  • RQ2: Can we detect asset management vulnerabilities in closed-source contracts? Even after deobfuscation, detecting such vulnerabilities can be challenging, given the complexity of MEV bots. Existing tools often timeout when applied to MEV bots, even without obfuscation.
@@
Challenges.  Numerous tools [15, 12, 28, 34, 30] are available to identify smart contract bugs. However, we highlight three limitations that render them inefficient for our tasks.
The first challenge is control-flow obfuscation. Contracts may employ indirect jumps that depend on runtime values, obscuring the execution path. Existing tools typically misclassify such blocks as unreachable [30, 32, 40] or raise errors when encountering symbolic jump destinations [12, 28, 34, 58].
@@
Secondly, detecting asset management vulnerabilities requires a fine-grained analysis of a smart contract’s logic, and simple pattern matching, as done by existing tools, does not suffice. E.g., JACKAL [34] flags CALL instructions with a fixed function selector as safe, overlooking vulnerabilities where function parameters can be manipulated. Mythril [12] only flags vulnerable CALL instructions when calldata can be completely controlled by the adversary, missing attacks that only require partially adversarial calldata.
The third challenge is the complexity of our real-world targets, which involve trading, flash loans, DEX swaps, and other interactions. This results in path explosion and expensive constraint solving, making symbolic and concolic execution inefficient without scalable methods for input generation.
Our Methods.  In this paper, we present the design and implementation of skanf  [^2] The name skanf stands for Sen, Kaihua, Aviv ’n Fan., a tool for detecting asset management vulnerabilities in closed-source and obfuscated smart contracts. skanf addresses the above challenges with the following three key ideas.
First, we de-obfuscate control flows by rebuilding jump tables. We observe that, unlike x86-64 or ARM64, where jumps can target any address, valid jump destinations in EVM (Ethereum Virtual Machine) must be marked with JUMPDEST [75]. This enables us to reconstruct the control flow and instrument the bytecode with a branch table to improve the coverage of subsequent analyses.
Second, to detect asset management vulnerabilities, we combine symbolic execution and taint analysis to identify vulnerable calls that adversaries can trigger with malicious parameters, including those where only critical parameters are adversary-controlled.
Third, we use concolic execution to speed up symbolic execution and mitigate path explosion through concrete seed inputs [61]. To obtain high-quality seed inputs at a large scale, we observe that historical transactions, when available, can be leveraged to automatically extract input seeds to improve both efficiency and detection efficacy. In the absence of historical transactions, our symbolic execution still functions correctly, though without the efficiency gains from seed inputs.
@@
To summarize, we make the following contributions:
@@
  • Building on these two techniques, we implement skanf, which takes the bytecode and the historical transactions of a given smart contract as input and outputs exploits if the smart contract is vulnerable. skanf can be used defensively by developers to detect and fix vulnerabilities before deployment or interaction.
@@
  • We evaluate the efficacy of skanf against real-world closed-source and obfuscated smart contracts. skanf identifies 1,030 vulnerable contracts from 6,554 MEV bots and automatically exploits 394 of them. The potential losses of these vulnerabilities exceed \$10.6M.
@@
  • We discover 104 real-world attacks targeting real-world contracts, which we categorize as MEV phishing attacks. These have caused an estimated loss of \$2.76M. Of the attacked smart contracts, 28 are flagged by skanf.
2 Preliminaries
2.1 Ethereum and Smart Contracts
@@
Ethereum.  Ethereum [75] is the second-most popular blockchain, and the foremost smart contract-enabled blockchain. It relies on a Proof-of-Stake (PoS) mechanism, wherein actors can become part of the set of validators operating the blockchain by staking (i.e., depositing) Ethereum’s native token, ETH. Ethereum’s mechanism sequentially assigns one validator to propose a block, i.e., a batch of transactions, every 12 seconds. The validator chosen to act as a proposer has a free hand in constructing its block.
@@
MEV supply chain.  Today’s Ethereum block construction pipeline involves several specialized entities. At one end, we have searchers who identify profitable opportunities, e.g., transactions that can be front-run [90], and assemble transaction bundles that exploit them. These bundles are then sent to builders, who specialize in constructing profitable blocks, and who compete in an auction-esque process to have their blocks chosen by upcoming proposers.
@@
Smart contract.  Smart contracts are programs stored and executed on the blockchain. They can be invoked through transactions. On Ethereum, contracts are deployed as Ethereum Virtual Machine (EVM) bytecode, typically written in high-level languages such as Solidity [68] and compiled for deployment [75]. Developers may optimize the contracts through compiler settings [69] or by writing bytecode directly.
Executed in the EVM, contracts are sequences of opcodes such as CALL for external interactions and JUMP/JUMPI for unconditional and conditional branching. The EVM is a stack-based environment where operations manipulate a last-in-first-out (LIFO) stack. Although contracts are stored on-chain as bytecode, their source code and Application Binary Interface (ABI) can be voluntarily published on platforms such as Etherscan [21]; without the ABI, it is difficult to determine how to interact with the contract.
@@
Indirect jump.  An indirect jump occurs when a jump instruction (JUMP/JUMPI) is executed and the top stack value (i.e., the jump destination) is not a statically known constant, but rather a runtime-computed value. Obfuscation techniques often rely on indirect jumps to hide control flow, making it difficult to statically determine which path may be executed.
@@
External call.  A smart contract can invoke another contract using one of four opcodes: CALL, CALLCODE, DELEGATECALL, and STATICCALL. These opcodes enable the caller to pass along Ether (value), execution gas, and input data (calldata), facilitating inter-contract interactions. They differ in how the execution context is propagated, particularly with respect to msg.sender, tx.origin, and storage access. In this paper, we focus on the CALL instruction, which underlies common patterns such as calling an ERC-20 token’s transfer function; we refer to such invocations as external calls.
@@
We refer to the contract that an external call invokes as the target address. The calldata consists of two parts: the first four bytes, the function selector, which identify the target function in the target contract; and the remaining bytes, which specify the function arguments.
@@
Asset management.  In this paper, we focus on ERC-20 token assets held by smart contracts. The ERC-20 standard [19] supports two transfer mechanisms: an account can call transfer to send tokens directly, or use approve to authorize another account to move tokens with transferFrom.
2.2 Smart Contract Obfuscation
Several obfuscation techniques have been proposed, including layout obfuscation [11, 84], data flow obfuscation [11, 84], control flow obfuscation [11, 45], and preventive transformations [11]. We focus on tackling control flow obfuscation because it prevents existing detection tools from functioning correctly, and it is commonly present among MEV bots (see [45] for empirical evidence).
We confirmed that the layout and data flow obfuscation do not interfere with existing analysis tools. We use the state-of-the-art obfuscator, BiAn [84] to apply layout and data flow obfuscation to 67 vulnerable smart contracts collected from the Smart Contract Weakness Classification registry [65]. Mythril can still detect vulnerabilities in all of the obfuscated contracts. In the interest of space, we relegate details to App. C. Preventive transformations, such as self-destructing after execution to make bytecode unavailable, are also less concerning due to blockchain’s transparency and immutability; even a self-destructed smart contract’s deployment transaction remains permanently recorded and publicly available for analysis.
2.3 Threat Model